# Information theory for data-driven model reduction in physics and biology

### v3

@@ I Model reduction as a compression problem
is known as the information bottleneck (IB) [38, 39]. It (and related variants [42, 43]) has been used in a variety of contexts, from document clustering [42] to neural codes [44] to (most relevant for us) dynamical systems [36, 45, 46], as well as in deep-learning contexts [47, 41, 48].
The IB objective ((3)) is a functional over stochastic functions, which may be represented by a conditional probability distribution $h(x_{t})\sim p(h|x_{t})$ (we denote random variables by uppercase $X_{t}$ and their values by lowercase $x_{t}$). Crucially, the parameter $\beta$ controls the trade-off between compression and prediction. For small $\beta$ the compression term dominates and the optimal encoder is trivial, losing all information about the system. For intermediate $\beta$ the compression term does not allow $X_{t}$ to be completely represented by $H_{t}$, so features of $X_{t}$ must compete to pass through to the encoding variable (Fig. 1b). These selected features are reflected in the IB encoder
$p^{*}(h_{t}|x_{t})=\text{arg min}\;\mathcal{L}_{\text{IB}}$ (4)
which provides the optimal trade-off between retained information and predictability [37]. In the optimal encoder, information about $X_{t}$ has been removed both by reducing the dimensionality, as well as by injecting noise into $h_{t}$. Increasing $\beta$ can impact both of these, either by increasing the dimensionality to learn new features, or by reducing noise to better identify the features that have already been found. In order to “perfectly” identify the features given a fixed set of components, in the second step we modify the IB encoder by discarding the part attributable to noise alone. This effectively increases the coupling of the reduced variables to the already-learned features. We shall see that this procedure isolates deterministic, interpretable reduced variables in both discrete and continuous settings.
Figure 1: Interpretable dynamical variables in reduced models via the information bottleneck. (a) In general model reduction aims to both learn an encoding and reduced dynamics, so that the reduced variable serves as a useful proxy that may be decoded for an estimate of the full system. Requiring the reduced variable to only be informative of the full state decouples these two steps (see SI Section 2). (b) The information bottleneck compresses high-dimensional state variables $x_{t}$, into simpler encoding variables $h_{t}$ with a controllable trade-off between the degree of compression and the predictive power about the system’s future. With deep neural networks, the encoding can be computed directly from data of observed fluid flows (left) or biological datasets, such as fluorescently labeled bacteria colonies (right). In general, the state of the variable $x_{t}$ may comprise time-lagged variables of the intensity field, $x_{t}=\{I_{t},I_{t+\Delta t}\}$ (right). The amount of compression is determined by the “width” of the bottleneck $\beta$ (see (3)). The resulting compressed, or encoded, variables $h_{t}$ represent collective variables most predictive of the system’s future. (c) Schematic evolution of the encoder $p(h_{t}|x_{t})$ for varying compression strength $\beta$. For low $\beta$ (high compression), the encoder is trivial and forgets everything about the input $x_{t}$. After the first IB transition at $\beta_{1}$, the encoder becomes non-trivial by gaining some dependence on $x_{t}$; some features of the input are able to pass through the bottleneck. At subsequent IB transitions, additional features are learned. (d) The point spectrum of the transfer operator contains several slowly decaying modes (red). We show that the most predictive variables that IB systematically extracts correspond to the slowest eigenfunctions of the transfer operator, associated to eigenvalues $\Lambda_{i}$ with $|\Lambda_{i}|\approx 1$. In fluid flows, the slowest-decaying eigenfunctions typically represent large-scale coherent patterns of the flow field, while faster-decaying eigenfunctions correspond to variations over shorter length scales. In general the system may exhibit an essential spectrum, however for noisy systems this is small (see main text).
II The optimal encoder in terms of the transfer operator
In any realistic experimental setting, the presence of noise or uncertainty means we cannot predict precisely the future state of a system but instead only a likely distribution of future states. Our prediction of the state at $\Delta t$ in the future is then represented mathematically as $p(x_{t+\Delta t}|x_{t})$, the probability of observing state $x_{t+\Delta t}$ given the current state $x_{t}$. For Markovian, or “memoryless” dynamics, this conditional probability distribution completely characterizes the dynamics of the system, and determines how probability distributions evolve in time. In this work we assume that the systems we consider are (or can be made) Markovian; for a detailed discussion of how to measure the memory and how it changes our below results, see SI Sections 3 and 4D. The evolution of probability distributions can then be understood as the action of a linear transfer operator $U^{\Delta t}$ on probability distributions:
$p_{X_{t+\Delta t}}(x_{t+\Delta t})=\int p(x_{t+\Delta t}|x_{t})p_{X_{t}}(x_{t})\text{d}x_{t}\equiv U^{\Delta t}[p_{X_{t}}]$ (5)
$U^{\Delta t}$ can be decomposed as follows (using standard notation from quantum mechanics and explained in the SI):
@@ II The optimal encoder in terms of the transfer operator
Our key observation is that the optimal IB encoder in (4), which we will use as a “filter” to extract relevant features, can be expressed in terms of the eigenvalues $\lambda_{n}$ and left eigenfunctions $\phi_{n}$ of (the generator of) $U$,
@@
where $f_{n}(h_{t})$ are factors that do not depend on $x_{t}$. For an outline of the mathematical steps leading to this see Methods, as well as SI Section 4. The combined effect of the factors in the exponent determines what features the encoder learns about the state $x_{t}$. In general, there may be a large number of non-zero factors $f_{n}$ so that the learned features are difficult to interpret in terms of individual eigenfunctions $\phi_{n}$. However, things become simple in the limit of small $\beta$, or high compression. When $\beta$ is small the encoder is trivial: $p(h_{t}|x_{t})=p(h_{t})$ so the value $h_{t}$ is assigned at random with no regard to the state of the system. No feature has been learned, and all factors $f_{n}$ are equal to zero. As $\beta$ is increased, the encoder undergoes a series of transitions at $\beta=\beta_{1}<\beta_{2}<\beta_{3}$... where new features are allowed to pass through the bottleneck (Fig. 1c) [51, 52, 53, 54]. The first transition happens at a finite value of $\beta_{1}$ when the first, most predictive, feature is learned.
Surprisingly, we find that at the first IB transition the vector of $f_{n}$ coefficients is dominated by a single term $f_{1}$.
$p^{*}_{\beta}(h_{t}|x_{t})\approx\frac{p^{*}_{\beta}(h_{t})}{\mathcal{N}(x_{t})}\,\exp\left(\beta\,\text{e}^{\lambda_{1}\Delta t}\phi_{1}(x_{t})f_{1}(h_{t})\right)$ (8)
This is our main mathematical result, which we derive by considering a perturbative expansion of the IB objective for small $f_{n}$. A proof of Eq. 8 with clearly specified technical assumptions may be found in Section 4 of the SI. One notable assumption underpinning this result, in addition to Markovianity, is that the system is in a statistical steady state, so that $x_{t}$ is sampled from the steady state distribution. When this assumption is violated, eigenfunctions of the transfer operator may cease to carry information about the system’s dynamics and instead its pseudospectrum may become relevant [55]; see SI Section 4D.
@@
The above result shows that in the limit of high compression the encoder’s dependence on $x_{t}$ is given by the first left eigenfunction $\phi_{1}(x_{t})$, which is the slowest-varying function of the state under dynamics given by $U$. Therefore, Eq. 8 makes precise the intuitive statement that slow features are the most relevant for predicting the future. Our analytical result, while applying only to the dominant eigenfunction, is valid for arbitrary, including non-Gaussian, variables.
@@
We further observe numerically that this picture holds true more generally: at successive IB transitions, the learned features correspond to successive modes of the transfer operator. This is consistent with the exact results known for the specific case of Gaussian IB, where the encoder learns successive eigenvectors of a matrix related to the covariance of the joint $X_{t},X_{t+\Delta t}$ distribution at each IB transition [53].
@@
Due to the smallness of $f_{1}$ at the transition, the optimal IB encoder couples only weakly to the eigenfunction $\phi_{1}$, so that the dynamics of $h_{t}$ will be dominated by noise. Given the maximal number of available bits in memory, or variables used in an analytical model, the IB solution will thus not make the best use of the resources (IB is optimal with respect to the amount of information retained, but not the size of its representation [43]). However, we can use the optimal IB encoding as a filter to identify $\phi_{1}$, and then “distill” it by discarding the noisy part of the encoding, effectively increasing the coupling between $h_{t}$ and $\phi_{1}(x_{t})$ while keeping the number of components fixed. In practice, this is done by taking the most likely value for the reduced state, $h_{t}=\arg\max_{h}p(h|x_{t})$ (see SI Section 4E). Similar procedures are performed in other probabilistic dimensionality reduction approaches, e.g. probabilistic principal component analysis [56]. There one finds a stochastic mapping from the full state $x$ to the reduced representation $h$, and then discards the noise to obtain the deterministic reduction.
@@
Together this lays the foundation of an operational prescription for model reduction which is built on IB for identifying relevant, slow, features. As we show later, this insight can be leveraged to systematically learn these slow variables directly from data with neural networks [41] which in turn may be used as inputs to an equation learning pipeline.
III Information decay and the spectrum of the transfer operator
Figure 2: Information loss of a Brownian particle in a double well potential. (a) Fixed point (FP) diagram of the dynamics given by Eq. 9 for zero noise. There is a bifurcation at $\mu=0$ where the stable FP at $x=0$ becomes unstable and two new stable FPs appear at $\pm\sqrt{\mu}$. (a,i) The corresponding potential $V(x)$ for varying $\mu$, with the emergence of a double-well structure for $\mu>0$. (a,ii) Dynamics of the system Eq. 9 for varying values of $\mu$ corresponding to the potentials above, with uniformly-distributed initial conditions. (a,iii) Loss of information between the initial condition and the future state. Inset shows scaling given by the first eigenvalue of the transfer operator. (b) Spectrum of the transfer operator $U$, showing a pile-up of eigenvalues for $\mu\gtrsim 0$. These are related to the eigenvalues of its infinitesimal generator by $\Lambda_{i}=\text{e}^{\lambda_{i}\Delta t}$. (c) Mutual information between the present and future state for varying time delay $\Delta t$ and bifurcation parameter $\mu$.
Figure 3: “Knowing when to stop”. The spectral properties of the transfer operator determine the necessary complexity (i.e. “when to stop” [33]) of the reduced model, which we show is also visible in information theoretic metrics. (a) Four-well potential in which a Brownian particle fluctuates. The magnitude $\sigma$ of the fluctuating noise is related to an energy scale $E_{\sigma}=\sigma^{2}$. (b) Information contained in the encoding variable $H_{t}$ about the future state $X_{t+\Delta t}$ for varying levels of noise and alphabet sizes $N_{H}$. (c) Information gain achieved by increasing the alphabet size by a single variable. This is the discrete derivative of the curve in (b). (d) Spectrum of the transfer operator for changing values of noise amplitude.
Figure 4: Non-Markovianity for a Brownian particle in a double well potential. (a) Conditional mutual information $I(H_{t-\Delta t},H_{t+\Delta t}|H_{t})$ as a function of $N_{H}$. (b) Two possible scenarios (left and right) with the same value of $h_{t}$, but different $h_{t-1}$. Here, the circles represent the value of the full state $x_{t}$ while the colored regions represent the value of the reduced state $h_{t}\in\{1,2\}$ which are mapped to for each $x_{t}$. (c) Illustration of the transition probability which takes into account the previous state $h_{t-1}$ (left), compared to the Markovian transition probability obtained by marginalizing out $h_{t-1}$ (right). In a Markovian system, these transition probabilities would coincide. (d-e) Same as in (h), but for $N_{H}=4$ states.
@@ III Information decay and the spectrum of the transfer operator
To develop intuition for information in a dynamical system, we turn to the simple example of a Brownian particle in a double-well potential. This may represent, for example, a molecule with a single degree of freedom that transitions between two metastable configurations [57]. In the overdamped limit the state of the particle is completely determined by its position $X_{t}\in\mathbb{R}$, with dynamics given by the Langevin equation
@@
The deterministic dynamical system undergoes a bifurcation at $\mu=0$ (Fig. 2a). To quantify the amount of information about the future state $X_{t+\Delta t}$ contained in the initial state $X_{t}$ we compute their mutual information (see SI for details). The dynamics of $X_{t}$ are Markovian, so that for any sequence of times $t_{0}<t_{1}<t_{2}$, $p(X_{t_{2}}|X_{t_{1}},X_{t_{0}})=p(X_{t_{2}}|X_{t_{1}})$. From the data processing inequality, one has [58]
@@
where expectations are taken over the steady state distribution (see SI Section 3). Asymptotically, the information decay is set by the value of $\lambda_{1}$, the rate of decay of the slowest-varying function $\phi_{1}(x)$ under the dynamics of $U$. In the limit of infinite time, for any value of $\mu$ even weak noise will cause the mutual information to become zero as there is a non-zero (perhaps exponentially small) probability of hopping between the wells [49].
@@
The loss of information in time depends on the bifurcation parameter $\mu$ as summarized in Fig. 2c. Note the peak in $I(X_{t},X_{t+\Delta t})$ for small, positive $\mu$. This corresponds to dynamics where observation of $X_{t}$ strongly informs the future state; recall that the mutual information is maximized when the conditional entropy $\mathcal{H}(X_{t+\Delta t}|X_{t})\approx 0$ (see SI Section 1). In contrast, for large positive or negative $\mu$, $X_{t}$ is not as informative of $X_{t+\Delta t}$ even for small times: the initial state is quickly forgotten as the particle approaches the bottom of the single (for $\mu<0$) or double (for $\mu>0$) well.
@@
This phenomenon is reminiscent of critical slowing down, which occurs in the noise-free system as $\mu$ passes through the bifurcation at $\mu=0$. For the deterministic dynamics, the slowing down is reflected in the spectrum as a “pile up” of eigenvalues to form a continuous spectrum [49]. In the presence of noise, although the continuous spectrum becomes discrete [49, 50] there is still a pile-up of eigenvalues characterized by several eigenvalues becoming close to 1 (Fig. 2b). This pile-up gives rise to the information peak seen in Fig. 2c. The peak is not solely due to the closing spectral gap $\lambda_{1}-\lambda_{2}$, but is also impacted by the subdominant eigenvalues which accumulate at $\mu\approx 0.2$ (SI Fig. S2).
IV Knowing when to stop

### v4

@@ II.2.5 Continuous information-theoretic sizes help optimization
is known as the information bottleneck (IB) [41, 42]. It has been used (along with related mathematical [61, 51, 62] and computational [63, 47, 64] extensions) in a variety of contexts, from document clustering [61] to analysis of neural networks [65] to biological signaling networks [66, 67] to, most relevant for us, dynamical systems [39, 68, 69, 70, 71].
@@
II.2.6 Distilling the encoder yields relevant variables
The form of the optimal encoder depends strongly on the value of the trade-off parameter $\beta$. For small $\beta$ the compression term dominates and the optimal encoder is trivial, losing all information about the system. For intermediate $\beta$ the compression term does not allow $X_{t}$ to be completely represented by $H_{t}$, so features of $X_{t}$ must compete to pass through to the encoding variable (Fig. 1b). These selected features are reflected in an encoder that provides the optimal trade-off between retained information and predictability [72] (Appendix C). In the optimal encoder, information about $X_{t}$ has been removed both by reducing the dimensionality, as well as by injecting noise into $h_{t}$. Increasing $\beta$ can impact both of these, either by increasing the dimensionality to learn new features, or by reducing noise to better identify the features that have already been found.
We obtain the most relevant dimensionally-reduced variables with a two-step procedure. First, we minimize the IB objective to find the optimal encoder $p^{*}(h_{t}|x_{t})$. In order to best identify the features given a fixed set of components, in the second step we modify the IB encoder by discarding the part attributable to noise alone to recover a deterministic function $h(X_{t})$. This effectively increases the coupling of the reduced variables to the already-learned features. This “distillation” procedure is detailed in Sec. S4.8. We shall see that this procedure isolates deterministic, interpretable reduced variables in both discrete and continuous settings.
Figure 1: Data-driven identification of interpretable dynamical variables via the information bottleneck. (a) A reduced description of a system requires both an encoding procedure that isolates relevant variables, and a set of dynamical equations that evolve the relevant variables in time. In our information theoretic framework these two steps decouple (Appendix S2). (b) The information bottleneck compresses high-dimensional state variables $x_{t}$, into simpler encoding variables $h_{t}$ with a controllable trade-off between the degree of compression and the predictive power about the system’s future. With deep neural networks, the encoding can be computed directly from data of observed fluid flows (left) or biological datasets, such as fluorescently labeled bacteria colonies (right). In general, the state of the variable $x_{t}$ may comprise time-lagged variables of the intensity field, $x_{t}=\{I_{t},I_{t+\Delta t}\}$ (right). The amount of compression is determined by the “width” of the bottleneck $\beta$ (see Eq. (11)). The resulting compressed, or encoded, variables $h_{t}$ represent collective variables most predictive of the system’s future. (c) Schematic evolution of the encoder $p(h_{t}|x_{t})$ for varying compression strength $\beta$. For low $\beta$ (high compression), the encoder is trivial and forgets everything about the input $x_{t}$. After the first IB transition at $\beta_{1}$, the encoder becomes non-trivial by gaining some dependence on $x_{t}$; some features of the input are able to pass through the bottleneck. At subsequent IB transitions, additional features are learned. (d) The point spectrum of the transfer operator contains several slowly decaying modes (red). We show that the most predictive variables that IB systematically extracts correspond to the slowest eigenfunctions of the transfer operator, associated to eigenvalues $\Lambda_{i}$ with $|\Lambda_{i}|\approx 1$. In fluid flows, the slowest-decaying eigenfunctions typically represent large-scale coherent patterns of the flow field, while faster-decaying eigenfunctions correspond to variations over shorter length scales. In general the system may exhibit an essential spectrum, however for noisy systems this is small (see main text).
III The optimal encoder in terms of the transfer operator
III.1 General results
For Markovian (“memoryless”) dynamics, the conditional probability distribution $p(x_{t+\Delta t}|x_{t})\equiv p_{X_{t+\Delta t}|X_{t}}(x_{t+\Delta t}|x_{t})$ of observing state $x_{t+\Delta t}$ at any $\Delta t$ in the future given the current state $x_{t}$ completely characterizes the dynamics of the system, and determines how probability distributions evolve in time. The evolution of probability distributions can then be cast as
$p_{X_{t+\Delta t}}=U^{\Delta t}[p_{X_{t}}],$ (13)
the action on probability distributions $p_{X_{t}}$ of the linear transfer operator $U^{\Delta t}$ defined by
$U^{\Delta t}[p_{X_{t}}](x)\equiv\int p(x|x^{\prime})p_{X_{t}}(x^{\prime})\text{d}x^{\prime}$ (14)
The transfer operator $U^{\Delta t}$ can be decomposed as
[…]
Our key observation is that the optimal IB encoder in Eq. (12), which “filters” out the relevant features of $X_{t}$, can be expressed in terms of the eigenvalues $\lambda_{n}$ and left eigenfunctions $\phi_{n}$ of (the generator of) $U$,
@@ III.1 General results
where $f_{n}(h_{t})$ are factors that do not depend on $x_{t}$. For an outline of the mathematical steps leading to this see Appendix D and SI Section S4. The combined effect of the factors in the exponent determines what features the encoder learns about the state $X_{t}$. In general, there may be a large number of non-zero factors $f_{n}$ so that the learned features are difficult to interpret in terms of individual eigenfunctions $\phi_{n}$. However, things become simple in the limit of small $\beta$, or high compression. When $\beta$ is small the encoder is trivial: $p(h_{t}|x_{t})=p(h_{t})$ so the value $h_{t}$ is assigned at random with no regard to the state of the system. No feature has been learned, and all factors $f_{n}$ are equal to zero. As $\beta$ is increased, the encoder undergoes a series of transitions at $\beta=\beta_{1}<\beta_{2}<\beta_{3}$... where new features are allowed to pass through the bottleneck (Fig. 1c) [75, 76, 77, 78]. The first transition happens at a finite value of $\beta_{1}$ when the first, most predictive, feature is learned.
Theorem Consider a transfer operator $U$ with infinitesimal generator $\mathcal{L}_{U}$. Assume that $\mathcal{L}_{U}$ has a discrete spectrum with eigenvalues satisfying $0=\text{Re }\lambda_{0}>\text{Re }\lambda_{1}>\text{Re }\lambda_{2}\gg\text{Re }\lambda_{3}$.... Then, for $\beta$ just above the first IB transition $\beta_{1}$ the optimal encoder is given approximately by
@@
This result shows that in the limit of high compression, the encoder’s dependence on $x_{t}$ is given by the first left eigenfunction $\phi_{1}(x_{t})$, which is the slowest-varying function of the state under dynamics given by $U$. Therefore, Eq. (17) makes precise the intuitive statement that slow features are the most relevant for predicting the future. Our analytical result is valid for arbitrary, including non-Gaussian, variables.
Proof – We sketch here the main steps; see Appendix D and SI Section S4 for a more detailed treatment. To show Eq. (17) we perform a perturbative analysis of the full optimal encoder Eq. (16). Before the transition, the coefficients $f_{n}(h)$ ($n>0$) are zero. The instability at the first IB transition at $\beta=\beta_{1}$ corresponds to the emergence of negative eigenvalues in the Hessian of the IB loss which may be computed explicitly (Appendix D and SI Section S4). The values of $f_{i}$ can be read off from the eigenvector corresponding to the first negative eigenvalue, which show dependence only on $f_{1}$. The transition is continuous, so that after the transition the encoder depends only on $\phi_{1}(x)$ to the desired order. We verify our theorem numerically in the SI.
We have assumed that the systems we consider are (or can be made) Markovian; in the SI we provide a detailed discussion of how to measure the memory (Section S3 and SI Fig. S4) and how it changes our below results (Section S4.7 and SI Fig. S7). One notable assumption underpinning this result, in addition to Markovianity, is that the system is in a statistical steady state, so that $x_{t}$ is sampled from the steady state distribution. When this assumption is violated, eigenfunctions of the transfer operator may cease to carry information about the system’s long-time dynamics and instead its pseudospectrum may become relevant [79] (SI Section S4.7).
@@
We further observe numerically that this picture holds true more generally: at successive IB transitions, the learned features correspond to successive modes of the transfer operator (SI Figs. S5 and S6). This is reminiscent of the exact results known for the specific case of Gaussian IB, where the encoder learns successive eigenvectors of a matrix related to the covariance of the joint $X_{t},X_{t+\Delta t}$ distribution at each IB transition [77, 39].
@@
Due to the smallness of $f_{1}$ at the transition, the optimal IB encoder couples only weakly to the eigenfunction $\phi_{1}$, so that the dynamics of $h_{t}$ will be dominated by noise. Given the maximal number of available bits in memory, or variables used in an analytical model, the IB solution will thus not make the best use of the resources (IB is optimal with respect to the amount of information retained, but not the size of its representation [51]). However, we can use the optimal IB encoding as a filter to identify $\phi_{1}$, and then “distill” it by discarding the noisy part of the encoding, effectively increasing the coupling between $h_{t}$ and $\phi_{1}(x_{t})$ while keeping the number of components fixed. In practice, this is done by taking the most likely value for the reduced state, $h_{t}=\arg\max_{h}p(h|x_{t})$ (see SI Section S4.8). Similar procedures are performed in other probabilistic dimensionality reduction approaches, e.g. probabilistic principal component analysis [80]. There one finds a stochastic mapping from the full state $x$ to the reduced representation $h$, and then discards the noise to obtain the deterministic reduction.
@@
Together this lays the foundation of an operational prescription for model discovery which is built on IB for identifying relevant, slow, features. As we show later, this insight can be leveraged to systematically learn these slow variables directly from data with neural networks [47], which in turn may be used as inputs to an equation learning pipeline.
III.2 Information decay and the spectrum of the transfer operator: an example
@@
To develop intuition for information in a dynamical system, we turn to the simple example of a Brownian particle in a double-well potential. This may represent, for example, a molecule with a single degree of freedom that transitions between two metastable configurations [81]. In the overdamped limit, the state of the particle is completely determined by its position $X_{t}\in\mathbb{R}$, with dynamics given by the Langevin equation
@@
The deterministic dynamical system undergoes a bifurcation at $\mu=0$ (Fig. 2a). To quantify the amount of information about the future state $X_{t+\Delta t}$ contained in the initial state $X_{t}$ we compute their mutual information. The dynamics of $X_{t}$ are Markovian, so that for any sequence of times $t_{0}<t_{1}<t_{2}$, $p(X_{t_{2}}|X_{t_{1}},X_{t_{0}})=p(X_{t_{2}}|X_{t_{1}})$. From the data processing inequality, one has [82]
@@
where expectations are taken over the steady state distribution (SI Section S3). Asymptotically, the information decay is set by the value of $\lambda_{1}$, the rate of decay of the slowest-varying function $\phi_{1}(x)$ under the dynamics of $U$. In the limit of infinite time, for any value of $\mu$ even weak noise will cause the mutual information to become zero as there is a non-zero (perhaps exponentially small) probability of hopping between the wells [73].
Figure 2: Information and the spectrum of the transfer operator. (a) Fixed point (FP) diagram of the dynamics given by Eq. (18) for zero noise. There is a bifurcation at $\mu=0$ where the stable FP at $x=0$ becomes unstable and two new stable FPs appear at $\pm\sqrt{\mu}$. (a,i) The corresponding potential $V(x)$ for varying $\mu$, with the emergence of a double-well structure for $\mu>0$. (a,ii) Dynamics of the system Eq. (18) for varying values of $\mu$ corresponding to the potentials above, with uniformly-distributed initial conditions. (a,iii) Loss of information between the initial condition and the future state. Inset shows scaling given by the first eigenvalue of the transfer operator. (b) Spectrum of the transfer operator $U$, showing a pile-up of eigenvalues for $\mu\gtrsim 0$. These are related to the eigenvalues of its infinitesimal generator by $\Lambda_{i}=\text{e}^{\lambda_{i}\Delta t}$. (c) Mutual information between the present and future state for varying time delay $\Delta t$ and bifurcation parameter $\mu$.
@@
The loss of information in time depends on the bifurcation parameter $\mu$ as summarized in Fig. 2b. Note the peak in $I(X_{t},X_{t+\Delta t})$ for small, positive $\mu$. This corresponds to dynamics where observation of $X_{t}$ strongly informs the future state; recall that the mutual information is maximized when the conditional entropy $\mathcal{H}(X_{t+\Delta t}|X_{t})\approx 0$ (SI Section S1). In contrast, for large positive or negative $\mu$, $X_{t}$ is not as informative of $X_{t+\Delta t}$ even for small times: the initial state is quickly forgotten as the particle approaches the bottom of the single (for $\mu<0$) or double (for $\mu>0$) well.
@@
This phenomenon is reminiscent of critical slowing down, which occurs in the noise-free system as $\mu$ passes through the bifurcation at $\mu=0$. For the deterministic dynamics, the slowing down is reflected in the spectrum as a “pile up” of eigenvalues to form a continuous spectrum [73]. In the presence of noise, although the continuous spectrum becomes discrete [73, 74] there is still a pile-up of eigenvalues characterized by several eigenvalues becoming close to 1 (Fig. 2c). This pile-up gives rise to the information peak seen in Fig. 2b. The peak is not solely due to the closing spectral gap $\lambda_{1}-\lambda_{2}$, but is also impacted by the subdominant eigenvalues which accumulate at $\mu\approx 0.2$ (SI Fig. S3).
Figure 3: “Knowing when to stop”. The spectral properties of the transfer operator determine the necessary complexity (i.e. “when to stop” [56]) of the reduced model, which is reflected in mutual information. (a) Four-well potential in which a Brownian particle fluctuates. The magnitude $\sigma$ of the fluctuating noise is related to an energy scale $E_{\sigma}=\sigma^{2}$. (b) Information contained in the encoding variable $H_{t}$ about the future state $X_{t+\Delta t}$ for varying levels of noise and alphabet sizes $N_{H}$. (c) Information gain achieved by increasing the alphabet size by a single variable. This is the discrete derivative of the curve in (b). (d) Spectrum of the transfer operator for changing values of noise amplitude.
IV Characterizing latent variables
In the next two sections, we focus on the case of a discrete latent space (like in Sec. III.2) in order to precisely characterize and demonstrate numerically two features of the latent variables: (i) the size of the latent space effectively satisfies a law of diminishing returns, in the sense that the information gain per additional extra state quickly decreases after some point, and (ii) the dynamics of latent variables are in general not Markovian, even if the underlying dynamics are Markovian; this non-Markovianity can be quantified and in some cases reduced.
IV.1 Knowing when to stop